# Neuro-Symbolic Embedding for Short and Effective Feature Selection via Autoregressive Generation

### v1

@@ 4.2.2. A Study of Integrating Variational Methods with Transformer
This experiment aims to answer: why do we use the benefits of using variational transformer instead of other encoder-decoder configurations (e.g., LSTM, VAE, transformer)? We evaluate our approach using three different encoder-decoder configurations. Initially, we employed LSTM as our sequential model, represented by ${}^{{}^{\prime}}$, and Figure LABEL:fig:configuration demonstrates that our generative feature selection framework outperforms the baselines. Subsequently, we replaced LSTM with a transformer, denoted as FSNS , and observed that FSNS yielded better performance compared to ${}^{{}^{\prime}}$. Finally, we compared our approach (denoted by FSNS) of integrating variational autoencoder into Transformer against a standalone transformer. Figure LABEL:fig:configuration shows the integration of variational learning into the Transformer consistently achieves higher accuracy. This can be attributed to the fact that the variational transformer can more effectively estimate and smooth the feature subset embedding space, resulting in improved feature subset generation.
@@ 4.2.3. On Effectiveness and Efficiency of Unsupervised Feature Selection
This experiment aims to answer: Does our method perform well even without supervision? We investigated the difference between unsupervised and supervised in terms of model performance. Figure LABEL:exp:u1 shows that our proposed is a generic framework that can achieve excellent results even without supervision signals. We use FSNS- to denote the unsupervised version of our method. Feature selection guided by an unsupervised Laplacian score based metric leads to an accuracy that is similar to usiing instead of supervised downstream task accuracy. This indicates that our model can be generalized to not just supervised setting but also unsupervised setting. We further compared different utility metrics of unsupervised feature sets and examined their impacts on our framework. Specifically, FSNS+ is our framework variant version that adopts feature set mean redundancy as a feature set utility metric. Table 2 shows that compared with redundancy, the aggregated Laplacian score-guided model can achieve better results on most of the data sets.
Moreover, we study the time costs of training data collection when comparing unsupervised and supervised implementations. Figure LABEL:exp:u2 shows that the unsupervised model can save a significant amount of time on training data collection, especially when dealing with large datasets. This is because using unsupervised strategies can avoid the need to repeatedly testing downstream tasks to collect downstream task performance.
@@ 4.2.4. On Effectiveness of Minimizing Redundancy
This experiment aims to answer: How does the strategy of minimizing redundancy impact the quality and size of a selected feature subset? Firstly, we studied the difference before and after applying the strategy of minimizing redundancy. Table 3 shows the results before and after redundancy minimization. Specifically, FSNSM denotes the redundancy minimized model using mutual information. Compared to FSNS, FSNSM can better reduce the redundancy of feature subsets with a minimal loss in accuracy. We examined the effect of using number of features as feature redundancy, as shown in Table 4. Overall, all three methods are effective in reducing the number of features, especially in noisy datasets like openml-586. Another interesting finding is that by minimizing feature-feature redundancy, we can jointly improve feature subset performance and decrease feature sizes.
Table 2. Comparison of different unsupervised performance evaluation methods
@@
Table 3. Results before and after redundancy minimization. We compare from downstream task accuracy and feature subset redundancy.
@@
This experiment aims to answer: How does our RL data collector and data augmentation method impact model performance? We conducted two control group experiments on three distinct datasets to investigate the roles of the data collector and data augmentation. Firstly, we compared the reinforcement learning-based data collector with random selection. Figure LABEL:exp:data_collection shows that random selection significantly decreased performance, even falling below baseline performance levels in several instances. This can be explained by the fact that the reinforcement learning-based data collector, in comparison to random selection, has the capacity to explore high-quality and diverse training data. Consequently, it constructs a more effective embedding space to advance the search of optimal feature subsets. Moreover, we compared the performances with data augmentation with the performances without data augmentation on the three datasets. Figure LABEL:exp:data_augmentation shows that the x-axis represents the number of times that we randomly shuffled the collected training set, and the y-axis represents the performance improvement percentages in a downstream task. We consistently observed that the downstream task performance improved as the number of feature shuffle operations increased. This indicates shuffling-based augmentation can diversify training data, overcome sparsity, and boost the effectiveness of generated feature subsets.
@@ 4.2.6. Robustness Check
This experiment aims to answer: Is our method robust when different ML models are used as a downstream task? Feature selection methods depend on evaluating the currently selected feature subset at each step with a downstream predictive model to identify the optimal feature subset. In our design, we employed the Random Forest for evaluation due to its stability and robustness. However, more rigorously, a feature selection method should be able to generalize to diverse downstream models and the selected features can achieve robust performances over diverse downstream models. For this purpose, we tested all the best feature subsets selected by baselines and our method with 5 widely used downstream predictive models, including support vector machine (SVM), XGBoost (XGB), K-nearest neighborhood (KNN), and decision tree (DT) with the SVMguide3 dataset in terms of the F1 score. Table 5 shows that, regardless of the downstream model used to evaluate the optimal feature subset generated by FSNS, the F1 score outperforms that of the baseline and the original feature set. Therefore, our method exhibits downstream model-agnostic robustness and reliability.
Table 5. The impact of different downstream models on the results of the SVMGuide3 dataset.
@@ 4.2.7. Examining the Scalability with Respect to Feature Size
This experiment aims to answer: Is the time and space complexity of our model relatively stable as the number of features increase? In other words, can our method efficiently handle large feature sets while maintaining acceptable running time and manageable parameter sizes? We selected 7 datasets with feature sizes ranging from 21 (the svmguide3 dataset) to 10,936 (ap_omentum_ovary dataset, referred to as ap_10936 in Figure LABEL:exp:complexity) to evaluate the training time, data collection time, search time, and needed parameter size with respect to feature number. Figure LABEL:exp:complexity shows that, as the feature size increases from 21 of svmguide3 to 10,936 of ap_omentum_ovary (470 times that of 21), The time spent on training, data collection, and searching increased by 20, 74, and 169 times, respectively. The parameter sizes increase 8 times compared to their original parameter size with 21 features. Our method demonstrates promising scalability.
@@
This experiment aims to answer: Can our model resist to noise and recognize high-quality features? We compared our method with an automated discrete choice model: MARLFS (Liu et al., 2019) on openml_607 and openml_618. Both datasets consist of 5 real features and 45 noise features. Figure LABEL:exp:case_study shows: 1) our method selected 4 informative features but only 3 noise features from on openml_607; 2) MARLFS selects 4 informative features but 23 noise features. Similarly, on the data set of openml_618, our method selects 4 informative features and MARLFS selects 5 informative features. But, MARLFS chose far more noise features than our method. Our method demonstrates the ability to effectively denoise features and identify high-quality features.
@@ 5.1. Feature Selection
Firstly, we review some classical works in feature selection. For instance, statistical methods including the Pearson correlation test (Meng et al., 1992), the student’s t-test (Owen, 1965), the Wilks’ lambda test (El Ouardighi et al., 2007), complementary and consensus learning (Huang et al., 2023) and the ANOVA F-test (St et al., 1989) are widely used in feature selection. These methods typically assess the importance of each feature from a statistical perspective, where higher scores indicate greater importance of the corresponding feature. In addition, the least absolute shrinkage and selection operator (LASSO) (Tibshirani, 1996) adds L1-regularization to simple linear regression, which can shrink the coefficients of unimportant variables to zero, thereby achieving feature selection. Decision tree algorithm has also been applied to feature selection (Sun and Hu, 2017), evaluating the importance of features by information entropy.
In recent years, reinforcement learning shows promising potential in feature learning (Wang et al., 2022; Xiao et al., 2023a, 2024). These methods (Kroon and Whiteson, 2009; Liu et al., 2023; Fard et al., 2013) create an agent to take actions in deciding the selection and deselection of all feautures. Each action yields a reward to assess its effectiveness, aiming to improve subsequent decisions on action. Furthermore, multi-agent reinforcement learning feature selection (MARLFS) is proposed (Liu et al., 2019, 2021) to employ an agent for each feature. Multi-agents interact with each other, collectively engaging in the search for the optimal feature subset.
@@
Due to its powerful feature representation capability and generalization, deep learning plays significant roles in various fields (Xiao et al., 2023b; Wang et al., 2024b). Some existing studies try to incorporate deep learning into feature selection, achieving competitive results. Li et al. (Li et al., 2016) develop a deep feature selection (DFS) model. Using elastic net, they introduce a sparse one-to-one linear layer between the input layer and the first hidden layer of a multi-layer perceptron. Then, the important features are selected according to their weights in the input layer. Also in (Shrikumar et al., 2017), a method called Deep Learning Important Features (DeepLIFT) is proposed to score features using a back propagation-like algorithm. Nezhad et al. (Nezhad et al., 2016, 2018) employ autoencoder to obtain high-level feature representations. Then, the features are selected by a classical feature selection method and applied to a downstream model. The number of nodes in the autoendoer hidden layers, is adjusted to continuously optimize the structure of the network based on the downstream model results.

### v2

@@ 4.2.2. A Study of Integrating Variational Methods with Transformer
This experiment aims to answer: why do we use the benefits of using variational transformer instead of other encoder-decoder configurations (e.g., LSTM, VAE, transformer)? We evaluate our approach using three different encoder-decoder configurations. Initially, we employed LSTM as our sequential model, represented by ${}^{{}^{\prime}}$, and Figure 3 demonstrates that our generative feature selection framework outperforms the baselines. Subsequently, we replaced LSTM with a transformer, denoted as FSNS , and observed that FSNS yielded better performance compared to ${}^{{}^{\prime}}$. Finally, we compared our approach (denoted by FSNS) of integrating variational autoencoder into Transformer against a standalone transformer. Figure 3 shows the integration of variational learning into the Transformer consistently achieves higher accuracy. This can be attributed to the fact that the variational transformer can more effectively estimate and smooth the feature subset embedding space, resulting in improved feature subset generation.
@@ 4.2.3. On Effectiveness and Efficiency of Unsupervised Feature Selection
This experiment aims to answer: Does our method perform well even without supervision? We investigated the difference between unsupervised and supervised in terms of model performance. Figure 4(a) shows that our proposed is a generic framework that can achieve excellent results even without supervision signals. We use FSNS- to denote the unsupervised version of our method. Feature selection guided by an unsupervised Laplacian score-based metric leads to an accuracy that is similar to using instead of supervised downstream task accuracy. The potential insight is that the Laplacian score can approximately evaluate feature importance, guiding the model to select significant features and eliminate redundant ones. This indicates that our model can be generalized to not just supervised setting but also unsupervised setting. We further compared different utility metrics of unsupervised feature sets and examined their impacts on our framework. Specifically, FSNS+ is our framework variant version that adopts feature set mean redundancy as a feature set utility metric. Table 2 shows that compared with redundancy, the aggregated Laplacian score-guided model can achieve better results on most of the datasets. When dealing with complex datasets, particularly noisy ones, the ability of redundancy as feedback to perform feature selection is rather limited. In contrast, the aggregated Laplacian score-guided model is more inclined to retain useful features while reducing the redundancy of the feature set.
Moreover, we study the time costs of training data collection when comparing unsupervised and supervised implementations. Figure 4(b) shows that the unsupervised model can save a significant amount of time on training data collection, especially when dealing with large datasets. This is because using unsupervised strategies can avoid the need to repeatedly test downstream tasks to collect downstream task performance.
(a) Downstream task accuracy
(b) Time cost of RL data collector
Figure 4. Results of supervised and unsupervised data collection methods. We compared (a) downstream task accuracy across four datasets, and (b) RL data collection time.
@@ 4.2.4. On Effectiveness of Minimizing Redundancy
This experiment aims to answer: How does the strategy of minimizing redundancy impact the quality and size of a selected feature subset? Firstly, we studied the differences before and after applying the strategy of minimizing redundancy. Table 3 shows the results before and after redundancy minimization. Specifically, FSNSM denotes the redundancy minimized model using mutual information. Compared to FSNS, FSNSM can better reduce the redundancy of feature subsets with a minimal loss in accuracy. This suggests that the dual-objective optimization approach endows the model with the ability to discern succinct and effective feature subsets, thereby enhancing the application on downstream tasks with limited computational resources. We further examined the effect of using the number of features as feature redundancy, as shown in Table 4. Overall, all three methods are effective in reducing the number of features, especially in noisy datasets like openml-586. Another interesting finding is that by minimizing feature-feature redundancy, we can jointly improve feature subset performance and decrease feature sizes. This indicates that in the presence of noise within the dataset, reducing feature redundancy not only decreases the number of features but also eliminates noise, thereby enhancing the model’s performance.
Table 2. Results of using different performance evaluation methods (i.e., downstream model performance, Laplacian score, and feature set mean redundancy. )
@@
Table 3. Results before and after redundancy minimization. We compare downstream task accuracy and feature subset redundancy.
@@
This experiment aims to answer: How does our RL data collector and data augmentation method impact model performance? We conducted two control group experiments on three distinct datasets to investigate the roles of the data collector and data augmentation. Firstly, we compared the reinforcement learning-based data collector with random selection. Figure 5(a) shows that random selection significantly decreased performance, even falling below baseline performance levels in several instances. This can be explained by the fact that the reinforcement learning-based data collector, in comparison to random selection, has the capacity to explore high-quality and diverse training data. Consequently, it constructs a more effective embedding space to advance the search of optimal feature subsets. Moreover, we compared the performances with data augmentation with the performances without data augmentation on the three datasets. Figure 5(b) shows that the x-axis represents the number of times that we randomly shuffled the collected training set, and the y-axis represents the performance improvement percentages in a downstream task. We consistently observed that the downstream task performance improved as the number of feature shuffle operations increased. This indicates that shuffling-based augmentation can diversify training data, overcome sparsity, and boost the effectiveness of generated feature subsets.
(a) Impact of RL data collector
(b) Impact of data augmentation
Figure 5. The effects of the RL data collector and data augmentation, (a) compare the downstream task accuracy achieved by random collection and the RL data collector, and (b) show the effect of different shuffling numbers on the accuracy
@@ 4.2.6. Robustness Check
This experiment aims to answer: Is our method robust when different ML models are used as a downstream task? Feature selection methods depend on evaluating the currently selected feature subset at each step with a downstream predictive model to identify the optimal feature subset. In our design, we employed the Random Forest for evaluation due to its stability and robustness. However, more rigorously, a feature selection method should be able to generalize to diverse downstream models and the selected features can achieve robust performances over diverse downstream models. For this purpose, we tested all the best feature subsets selected by baselines and our method with 5 widely used downstream predictive models, including support vector machine (SVM), XGBoost (XGB), K-nearest neighborhood (KNN), and decision tree (DT) with the SVMguide3 dataset in terms of the F1 score. Table 5 shows that, regardless of the downstream model used to evaluate the optimal feature subset generated by FSNS, the F1 score outperforms that of the baseline and the original feature set. Therefore, our method exhibits downstream model-agnostic robustness and reliability. The underlying driver is that we collect data corresponding to downstream models to train our model, formulate specific optimization strategies and achieve robust optimization search.
Table 5. Robustness check results. We compare model performance under different downstream models on the SVMGuide3 dataset.
@@ 4.2.7. Examining the Scalability with Respect to Feature Size
This experiment aims to answer: Is the time and space complexity of our model relatively stable as the number of features increases? In other words, can our method efficiently handle large feature sets while maintaining acceptable running time and manageable parameter sizes? We selected 7 datasets with feature sizes ranging from 21 (the svmguide3 dataset) to 10,936 (ap_omentum_ovary dataset, referred to as ap_10936 in Figure 6 to evaluate the training time, data collection time, search time, and needed parameter size with respect to feature number. Figure 6 shows that, as the feature size increases from 21 of svmguide3 to 10,936 of ap_omentum_ovary (470 times that of 21), The time spent on training, data collection, and searching increased by 20, 74, and 169 times, respectively. The parameter sizes increase 8 times compared to their original parameter size with 21 features. This is because, during the training phase, tokens are embedded into a low-dimensional space, making the time complexity and space complexity robust to the increase in feature dimensions. In the data collection phase, the time is primarily influenced by the computation of downstream model accuracy. However, the overall growth rate is acceptable, and our provided unsupervised collection method addresses this issue. Gradient-steered optimization search is generally time-efficient, even when the feature dimension is 10,936, a single search takes approximately 20 seconds, which has a minimal impact on the overall time complexity of the model. Therefore, our method demonstrates promising scalability.
(a) Training time
(b) Data collection time
(c) Search time
(d) Parameter size
Figure 6. Results of time and space complexity. We study the training time, data collection time, search time, and parameter size of FSNS as the number of features increase.
(a) Performance
(b) Redundancy
Figure 7. The hyparameter sensitivity analysis on SpamBase. The different colored lines denote different loss weight configurations, represented by [$\alpha,\beta,\delta$]. The horizontal axis represents the values of $\lambda$, while the vertical axis corresponds to performance and 1 - redundancy, respectively.
4.2.8. A Study of the Hyperparameter Sensitivity
This experiment aims to answer: Is the performance of FSNS sensitive to values of the trade-off parameters? We investigated the key hyperparameters, including the loss weight parameters $\alpha,\beta,\delta$, and the search trade-off parameter between performance and redundancy $\lambda$. We compared the performance and redundancy achieved by FSNS under different parameter configurations using the SpamBase dataset. The results are shown in Figure 7, where the different colored lines denote different loss weight configurations (represented as [$\alpha,\beta,\delta$]) and the horizontal axis denotes different $\lambda$ values. We can observe that different configurations of $\alpha,\beta$, and $\delta$ slightly affect the results with the same $\lambda$ value. This suggests that during the training phase, FSNS is robust to the hyperparameters and can effectively model the data distribution under various weight configurations. $\lambda$ is the trade-off parameter between performance and redundancy during the search phase. A higher $\lambda$ value makes the model more inclined to find feature subsets with lower redundancy. These findings provide insights into how the hyperparameter impacts the results of FSNS and how to choose the values of the hyperparameters to balance the performance and redundancy.
4.2.9. A study of the Traceability and Interpretability
This experiment aims to answer: Is the constructed feature space traceable and interpretable? According to the mutual information between features and label, we report in Figure 8 the top 10 essential features in the original and FSNS selected feature spaces respectively, with darker colors indicating more important features. We can observe that the valuable features (e.g., ”capital_run_length_longest”) are retained by FSNS, and the redundant features such as (e.g., ”word_freq_you”) are eliminated. The underlying driver is that long sequences of uppercase letters are often used in spam emails to grab attention and emphasize urgency. But the word ”you” serves a general communicative function. It is hard to distinguish spam emails according to the frequency of ”you”. In addition, new features (e.g., ”word_freq_all”) emerge among the top 10 features. This can be explained as “all” is frequently used to make broad promises or sweeping claims in spam emails. This result suggests that FSNS can explicitly trace the features through their names and interpret them through the semantic descriptions.
(a) Original feature space
(b) FSNS selected feature space
Figure 8. The top 10 important features identified in the original feature space and the FSNS selected feature space using the mutual information between features and label.
@@ 4.2.9. A study of the Traceability and Interpretability
This experiment aims to answer: Can our model resist noise and recognize high-quality features? We compared our method with an automated discrete choice model: MARLFS (Liu et al., 2019) on openml_607 and openml_618. Both datasets consist of 5 real features and 45 noise features. Figure 9 shows: 1) our method selected 4 informative features but only 3 noise features from on openml_607; 2) MARLFS selects 4 informative features but 23 noise features. Similarly, on the data set of openml_618, our method selects 4 informative features and MARLFS selects 5 informative features. But MARLFS chose far more noise features than our method. Our method demonstrates the ability to effectively denoise features and identify high-quality features.
(a) Openml_607
(b) Openml_618
Figure 9. Case study in terms of our method and MARLFS based on the two OpenML datasets. Each dataset consists of 5 real features and 45 noise features. FSNS & MARLFS: The number of feature subsets that our approach and RL-based method select from the OpenML dataset, respectively.
@@ 5.1. Feature Selection
Feature selection is crucial in improving the performance of downstream models (Peng et al., 2005; Wang et al., 2024a; Ying et al., 2024b, 2023, a; Hu et al., 2024; Ying et al., 2024c; Gong et al., 2024). Various classical works are widely used. For instance, statistical methods including the Pearson correlation test (Meng et al., 1992), the student’s t-test (Owen, 1965), the Wilks’ lambda test (El Ouardighi et al., 2007), complementary and consensus learning (Huang et al., 2023) and the ANOVA F-test (St et al., 1989). These methods typically assess the importance of each feature from a statistical perspective, where higher scores indicate greater importance of the corresponding feature. In addition, the least absolute shrinkage and selection operator (LASSO) (Tibshirani, 1996) adds L1-regularization to simple linear regression, which can shrink the coefficients of unimportant variables to zero, thereby achieving feature selection. Decision tree algorithm has also been applied to feature selection (Sun and Hu, 2017), evaluating the importance of features by information entropy.
In recent years, feature selection has been regarded as a discrete searching problem. Optimization algorithms such as metaheuristic algorithms (Sayed et al., 2019; Al-Sharaf et al., 2019; Zorarpacı and Özel, 2016) and reinforcement learning (Wang et al., 2022; Xiao et al., 2023a, 2024) show promising potential in this filed. Some metaheuristic algorithms including particle swarm optimization (PSO) (Ajibade et al., 2020), genetic algorithm (GA) (Oh et al., 2004), grasshopper optimization algorithm (GOA) (Al-Sharaf et al., 2019), Crow search algorithm (CSA) (Sayed et al., 2019), bee colony optimization (BCO) (Zorarpacı and Özel, 2016), and grey wolf optimizer (GWO) (Al-Tashi et al., 2019) have been applied to feature selection. These methods iteratively generate and evaluate new solutions. Reinforcement learning-based methods (Liu et al., 2019, 2023, 2021) create agents to take actions in deciding the selection and deselection of the features. Each action yields a reward to assess its effectiveness, aiming to improve subsequent decisions on action.
@@
Due to its powerful feature representation capability and generalizability, deep learning plays significant roles in various fields (Xiao et al., 2023b; Wang et al., 2024b). Some existing studies try to incorporate deep learning into feature selection, achieving competitive results. Li et al. (Li et al., 2016) develop a deep feature selection (DFS) model. Using elastic net, they introduce a sparse one-to-one linear layer between the input layer and the first hidden layer of a multi-layer perceptron. Then, the important features are selected according to their weights in the input layer. Also in (Shrikumar et al., 2017), a method called Deep Learning Important Features (DeepLIFT) is proposed to score features using a back propagation-like algorithm. Nezhad et al. (Nezhad et al., 2016, 2018) employ autoencoder to obtain high-level feature representations. Then, the features are selected by a classical feature selection method and applied to a downstream model. The number of nodes in the autoendoer hidden layers, is adjusted to continuously optimize the structure of the network based on the downstream model results.
@@ 6. Conclusion
Acknowledgment
This research was partially supported by the National Science Foundation (NSF) via the grant numbers: 2426340, 2416727, 2421864, 2421865, 2421803, and National academy of engineering Grainger Foundation Frontiers of Engineering Grants.